OpenStudy (anonymous):

sin θ cos 3θ + cos θ sin 3θ = 0

OpenStudy (campbell_st):

well it looks like \[\sin(\theta + 3 \theta) = \sin(4 \theta)\]

OpenStudy (campbell_st):

so I think you need to solve \[\sin(4 \theta) = 0\]